OpenStudy (anonymous):

50= (2x)^2/(1-x)(2-x)/3.. what is the final ans.???

sam (.sam.):

x=2.106942,0.9752496

OpenStudy (anonymous):

Assuming this is your question: \[\frac{\frac{(2x)^2}{(1-x)(2-x)}}{3}=50\] \[\frac{4x^2}{(1-x)(2-x)}=150\] \[4x^2=150(1-x)(2-x), x \neq 1, 2\] \[4x^2=150(2-3x+x^2)\] \[2x^2=75(x^2-3x+2)\] \[2x^2=75x^2-225x+150\] \[73x^2-225x+150=0\] \[x=\frac{225 \pm \sqrt{225^2-4(73)(150)}}{2(73)}=\frac{225 \pm 5\sqrt{273}}{146}\]